OpenStudy (anonymous):

the perimeter of a rectangle is 156 inches the length exceeds the width by 18 inches find the length and width find the lenght and width

OpenStudy (anonymous):

perimeter = 2x + 2y (where x and y are length and width respectively). thus x=y+18, so we have 156=2x+2y = 2(y+18) + 2y... solve for the value of y

OpenStudy (anonymous):

30 and 48

OpenStudy (anonymous):

once you have the value of y (the width) from the method i told you, then add 18 to get the length. and 30 and 48 are wrong...

OpenStudy (anonymous):

ignore that last part bward, i just brainfarted, those are definitely the right numbers